\begin{thebibliography}{9}

\bibitem{ahk} S.~Avramska-Lukarska, D.~Hundertmark,  H.~Kova\v r\'{\i}k: {\em Absence of positive eigenvalues for two-dimensional magnetic
    Schr\"odinger operators}, Calc. Var. Partial Differential Equations   62  (2023)  n. 63.

\bibitem{BK}   J.~Breuer,  H.~Kova\v r\'{\i}k: {\em Resonances at the Threshold for Pauli operators in Dimension Two},  Ann.~H.~Poincar\'e  25 (2024) 2839--2875.

\bibitem{cjn} H.~Cornean, A.~Jensen, G.~Nenciu: {\em Metastable States When the Fermi Golden Rule Constant Vanishes},  Commun. Math. Phys.  334 (2015) 1189--1218.

\bibitem{Ex85} P.~Exner: \emph{Open Quantum Systems and Feynman Integrals}; Fundamental Theories of Physics, vol.~6, D.~Reidel, Dordrecht 1985.

\bibitem{har} E.M.~Harrell: {\em Perturbation Theory and Atomic Resonances
Since Schr\"odinger’s Time},  in Spectral theory and mathematical physics. A festschrift in honor of Barry Simon’s 60th birthday. 
Proceedings of Symposia in Pure Mathematics 76, Pt. 1, 227-248 (2007).

\bibitem{Ho74} J.S.~Howland: {\em Puiseux series for resonances at an embedded eigenvalue}, Pacific J. Math. 86 (1974), 157--176.

\bibitem{hun} W.~Hunziker:  {\em Resonances, metastable states and exponential decay laws in perturbation theory}, Commun. Math. Phys. 132 (1990) 177--188.

\bibitem{J80} A. Jensen, \emph{Spectral properties of Schrödinger operators and time-decay of the wave functions. Results in $L^2(\bR^m)$, $m\geq5$}, Duke Math. J. 47 (1980), no. 1, 57--80.

\bibitem{JK} A. Jensen and T. Kato, \emph{Spectral properties of Schrödinger operators and time-decay of the wave functions}, Duke Math. J. 46 (1979), no. 3, 583--611.

\bibitem{JKmag} A. Jensen and H. Kova\v{r}\'{\i}k, \emph{Resolvent expansions of 3D magnetic Schr\"odinger operators and Pauli operators}, J. Math. Phys.  65  (2024), 19 pp.

\bibitem{JN} A. Jensen and G. Nenciu, \emph{The Fermi golden rule and its form at thresholds in odd dimensions}, Comm. Math. Phys. 261 (2006), no. 3, 693--727.

\bibitem{JN2} A. Jensen and G. Nenciu, \emph{Uniqueness results for transient dynamics of quantum systems}, Adventures in mathematical physics, 165--174, Contemp. Math., 447, Amer. Math. Soc., Providence, RI, 2007.

\bibitem{M} M. Murata, \emph{Asymptotic expansions in time for solutions of Schrödinger-type equations}, J. Funct. Anal. 49 (1982), no. 1, 10--56.

\bibitem{ort} A.~Orth: \emph{Quantum mechanical resonance and limiting absorption: the many body problem}, Commun. Math. Phys. 126 (1990), 559--573.


\bibitem{sw} A.~Soffer, M.~Weinstein: \emph{Time dependent resonance theory}, Geom.~Funct.~Anal. 8 (1988)  1086--1128.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{amsmath,amsthm,amsfonts,amsxtra,amssymb,upref}
\usepackage[mathscr]{eucal}
\usepackage[a4paper,margin=2.5cm]{geometry}

\usepackage{mathtools}
\usepackage{enumerate}
\usepackage{color}

\usepackage[parfill]{parskip}    % Activate to begin paragraphs with an empty line rather than an indent

\numberwithin{equation}{section}

\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]
\newtheorem{cor}[thm]{Corollary}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}
\newtheorem{assumption}[thm]{Assumption}

\newtheorem{rem}[thm]{Remark}

\newcommand{\eps}{\varepsilon}

\newcommand{\R}{\mathbb{R}}

\newcommand{\LL}{\mathcal{\big\langle}}
\newcommand{\RR}{\mathcal{\big\rangle}}

\newcommand{\cH}{\mathcal{H}}
\newcommand{\jap}[1]{\langle{#1}\rangle}

\DeclareMathOperator{\im}{Im}

\newcommand{\abs}[1]{\lvert{#1}\rvert}

\newcommand{\norm}[1]{\lVert{#1}\rVert}
\newcommand{\ket}[1]{\lvert{#1}\rangle}
\newcommand{\bra}[1]{\langle{#1}\rvert}
\newcommand{\ip}[2]{\langle{#1},{#2}\rangle}

\newcommand{\cO}{\mathcal{O}}

\DeclareMathOperator{\sign}{sign}

\newcommand{\bR}{\mathbf{R}}
\newcommand{\bC}{\mathbf{C}}
\newcommand{\bone}{\mathbf{1}}

\newcommand{\cB}{\mathcal{B}}
\newcommand{\cK}{\mathcal{K}}
\newcommand{\cU}{\mathcal{U}}
\newcommand{\We}{W_{\eps}}
\newcommand{\we}{w_{\eps}}
\newcommand{\wes}{w_{\eps}^{\ast}}
\newcommand{\Reff}{R_{\textrm{eff}}}
\newcommand{\Heff}{H_{\textrm{eff}}}
\newcommand{\tb}{\tilde{b}}
\newcommand{\Deps}{D_{\nu,\eps}}

\newcommand{\Esf}{\mathsf{E}}

\newcommand{\Ie}{I_{\eps}}
\newcommand{\xe}{x_0(\eps)}

\title{A modified Fermi Golden Rule at threshold for 3D magnetic Schrödinger operators}

\author{Pavel Exner\footnote{Doppler Institute for Mathematical Physics and Applied Mathematics, Prague  Czechia, and  Department of Theoretical Physics, Nuclear Physics Institute, Czech Academy of Sciences,
\v Re\v z near Prague, Czechia, \texttt{pavel.exner@ujf.cas.cz}}, \
Arne Jensen\footnote{Department of Mathematical Sciences, Aalborg
University, Thomas Manns Vej 23, DK-9220 Aalborg~\O{}, Denmark,
\texttt{matarne@math.aau.dk}} \ and  Hynek Kova\v{r}\'{\i}k\footnote{DICATAM, Sezione di Matematica, Università degli studi di Brescia, Via Branze 38, Brescia, 25123, Italy, \texttt{hynek.kovarik@unibs.it}}}

\begin{document}
\allowdisplaybreaks

\maketitle

\noindent {\bf Keywords:} resonances, threshold eigenvalue,  magnetic field, Fermi Golden Rule

\smallskip

\noindent {\bf MSC 2020:}  35Q40, 35P05, 81Q10

\begin{abstract}
In this paper we consider three-dimensional Schr\"odinger operators with a simple threshold eigenvalue. We show, under certain 
assumptions, that when a small magnetic field is introduced, this eigenvalue turns into a resonance 
in the time-dependent sense. We find the leading term in the asymptotic expansion of the imaginary part of the resonance and discuss 
the principal differences with respect to resonances induced by  weak electric fields obtained previously in the literature.  
\end{abstract}

\section{Introduction}

We consider the following operators on $\cH=L^2(\bR^3)$:
 %----------------%
\begin{align}
H_0&=-\Delta+V,\label{H0}\\
H_{\lambda}&=(P-\lambda A)^2+V,\label{H-lambda}
\end{align}
 %----------------%
where $P=-i\nabla$ and $V\colon\bR^3\to\bR$ is a bounded potential that decays sufficiently rapidly, say $\abs{V(x)}\leq C(1+\abs{x})^{-\beta}$ with $\beta>2$. The magnetic potential $A\colon \bR^3\to\bR^3$, which generates the magnetic field $B=\nabla \times A$, is also assumed to decay sufficiently rapidly, see Assumption \ref{assump}(2) below for details. Under these  assumptions  the point zero in the spectrum of $H_0$ may be an eigenvalue, a zero resonance, both, or neither; by zero resonance we mean that there exists a solution to $H_0\psi=0$ (in the sense of distributions), which satisfies  $\psi  \in L^2(\bR^d, \jap x^{ -\frac 12-\delta}\, dx)\setminus L^2(\bR^3)$ for any $\delta>0$.

In this paper we assume that zero, the essential spectrum threshold, is a simple eigenvalue of $H_0$, but not a resonance. 
Suppose that we turn on a magnetic field, i.e.~consider $H_{\lambda}$ for $\lambda$ small. The question that we address is the following: 

What happens to the eigenvalue embedded at the threshold zero? The expectation is that it either becomes an isolated eigenvalue close to zero or becomes a resonance. Notice that under our decay conditions on $V$ and $B$ the operator $H_\lambda$ cannot have an eigenvalue embedded in the continuous spectrum, \cite{ahk}. Assume that $\psi_0\in\cH$, $\norm{\psi_0}=1$, satisfies $H_0\psi_0=0$. We impose a condition that ensures the eigenvalue becomes a resonance, described by a metastable state. A time-dependent description of a resonance uses the survival probability of $e^{-itH_{\lambda}}\psi_0$ in the state $\psi_0$, cf.~\cite{Ex85, hun, JN, ort, sw}
and the review paper \cite{har}.
 We are going to show that
 %----------------%
\begin{equation}\label{exp-decay}
\ip{\psi_0}{e^{-itH_{\lambda}}\, \psi_0}=e^{-it(x_0(\lambda)-i\Gamma(\lambda))}+\cO(\lambda^{2p})
\quad\text{as $\:\lambda\to  0$} 
\end{equation}
 %----------------%
with the error term uniform in $t>0$. We have either $p=1/2$ or $p=2$, depending on properties of $H_0$ and $\psi_0$.
The point $x_0(\lambda)-i\Gamma(\lambda)$ with $\Gamma(\lambda)>0$ indicates the location of the resonance. Its imaginary part $\Gamma(\lambda)$, called the resonance width, is of main interest for us since it determines, up to the error term, the rate of decay of the survival probability $| \LL \psi_0, \, e^{-it H_\lambda}\, \psi_0 \RR|^2$, conventionally used to characterize the resonance lifetime.
We find the leading terms in the asymptotics of $x_0(\lambda)$ and $\Gamma(\lambda)$.

If an eigenvalue of $H_0$ is embedded in the continuum, $(0,\infty)$, then a perturbation $H_0+\lambda\widetilde{V}$ may turn it into a resonance the time evolution of which is described by \eqref{exp-decay}. In this case the imaginary part of the resonance position is typically given by the Fermi Golden Rule, $\Gamma(\lambda)=C_{\textup{FGR}}\lambda^2+\cO(\lambda^q)$ for some $q>2$, where the constant  $C_{\textup{FGR}}$ is generically positive. 

Our result gives a modification of the Fermi Golden Rule in the situation when the eigenvalue is situated at the continuous spectrum threshold, and the perturbation is given by a magnetic field.
In particular, our main result, Theorem \ref{main}, states that
 %----------------%
\begin{equation} \label{gamma-0}
\Gamma(\lambda) = \Gamma_0\, |\lambda|^3 +  \mathcal{O}(\lambda^4) \quad\text{as} \ \lambda\to 0.
\end{equation}
 %----------------%
The coefficient  $\Gamma_0$ is generically positive. However, Theorem \ref{main} provides an asymptotic expansion of $\Gamma(\lambda)$ also in the exceptional cases where $\Gamma_0=0$. 

The proof of our main result uses the factorization of the operator difference $H_\lambda-H_0$ recently developed in \cite{JKmag}. When combined with  the Schur-Lifschitz-Feshbach-Grushin (SLFG) formula, this factorization allows us to derive a threshold resolvent expansion of the operator $H_\lambda$ in the topology of weighted Sobolev spaces. Such a resolvent expansion then plays a crucial role in analyzing the location of the resonance, and hence in proving equation \eqref{gamma-0}.
 
It is illustrative to compare the behavior of the threshold resonances considered in this paper with those induced by a weak electric field, in other words arising from  a threshold eigenvalue as a result of an additive perturbation $\lambda W$. Such resonances were analyzed in \cite{JN}. There their existence was proved for $\lambda\to 0+$ under the condition $\LL \psi_0, W \psi_0\RR >0$, and it was shown that the leading term
of $\Gamma$ is generically of order $\lambda^{3/2}$, see \cite[Thm.~3.7]{JN}.
A comparison with equation \eqref{gamma-0} thus reveals two important differences. First, the quantity $\Gamma(\lambda)$ associated to resonances induced by a magnetic field is independent of the sign of $\lambda$. Second, its size is much smaller  with respect to the case of a perturbation by an electric field. This leads to a slower decay of the corresponding survival probability.

It should be also noticed that, by \cite[Prop.~1.1]{JN2},  a result of the type \eqref{exp-decay} defines the leading asymptotics in $x_0(\lambda)$ and $\Gamma(\lambda)$ uniquely, see Corollary \ref{cor-unique} for more details.

Some comments on the literature.
Recall that in other systems the presence of the unperturbed eigenvalue at the continuum edge may also result in the coupling constant power in the leading term different from two, cf. \cite[Example 2]{Ho74} or \cite[Example 3.2.5]{Ex85} for the Friedrichs model example; in the words of J.~Howland the gold from which the rule is made is then mixed with brass. Another notable feature of the Fermi Golden Rule is that the constant 
$C_{\textup{FGR}}$ might vanish. Such situations were studied in \cite{cjn}, where the subsequent term in the asymptotic expansion  of $\Gamma(\lambda)$ was calculated. Threshold resonances arising from a perturbation of 
the two-dimensional Pauli operator were found recently in \cite{BK}. In this case the asymptotic expansion of $\Gamma(\lambda)$ depends heavily on the flux of the magnetic field.

Our paper is organized as follows. In the next section we introduce the necessary notation and state the hypothesis on the magnetic and electric field. The essential resolvent expansions of the operator $H_\lambda$
are proved in Section \ref{sec-SLFG}, cf.~equation \eqref{eq-eff} and Lemmas \ref{lemma44}-\ref{lemma46}. The main result, Theorem \ref{main}, is stated and proved in Section \ref{sec-main}.

\section{Notation, definitions, and assumptions}
The basic Hilbert space is $\cH=L^2(\bR^3)$. The norm is denoted by $\norm{\cdot}$.
The weighted Sobolev spaces are denoted by $H^{k,s}(\bR^3)$, often written as $H^{k,s}$. The norm is denoted by $\norm{\cdot}_{k,s}$. See~\cite{JKmag} for further information on the weighted Sobolev spaces.
The bounded operators between two Hilbert spaces $\cK_1$ and $\cK_2$ are denoted by $\cB(\cK_1,\cK_2)$. In particular, we sometimes use the notation
\begin{equation*}
\cB(k_1,s_1;k_2,s_2)=\cB(H^{k_1,s_1},H^{k_2,s_2}).
\end{equation*}

We impose the following assumptions on $V$ and $A$ in the operators given by \eqref{H0} and \eqref{Heps}.

\smallskip

\begin{assumption}\label{assump}
\
\begin{enumerate}[(1)]
\item \label{assump1}
Assume $V\colon \bR^3\to\bR$ satisfies
\begin{equation}
\abs{V(x)}\lesssim\jap{x}^{-\beta_1},\quad \beta_1>2.
\end{equation}
\item \label{assump2}
Assume $B\in C^1(\bR^3;\bR^3)$ such that $\nabla\cdot B=0$. Assume $B=\nabla\times A$, $A\in C^2(\bR^3;\bR^3)$,  such that
\begin{equation}\label{decay}
\abs{B(x)}\lesssim \jap{x}^{-\beta_2-1},\quad
\abs{A(x)}\lesssim \jap{x}^{-\beta_2}, \quad \beta_2>2.
\end{equation}
\item \label{assump3}
Assume that $0$ is a simple eigenvalue of $H_0$, and furthermore that $H_0$ does not have a zero resonance. Let $\psi_0\in H^{2,0}$, such that $H_0\psi_0=0$ and $\norm{\psi_0}=1$. Assume $\psi_0$ is chosen to be \emph{real-valued}.
\end{enumerate}
\end{assumption}

\medskip

\begin{rem} 
For examples of operators $H_0$ satisfying condition (\ref{assump3}) we refer to \cite[Rem.~5.2]{JN} and \cite[Rem.~6.6]{JK}.
If $B$ satisfies the decay condition in \eqref{decay}, a construction of $A$ satisfying the corresponding decay condition is given in \cite{JKmag}.
\end{rem}

In the sequel it will be convenient  to work with the operator
\begin{equation}\label{Heps}
H_{\eps}=(P-\sqrt{\eps}A)^2+V, \qquad \eps>0.
\end{equation}
Our main result then will follow from a corresponding result for $H_\eps$ after identifying $\eps$ with $\lambda^2$, see the proof of Theorem \ref{main}.

We let
\begin{equation}
\We=H_{\eps}-H_0=\sum_{j=1}^3(-\eps^{1/2}P_jA_j-\eps^{1/2}A_jP_j)
+\eps\abs{A}^2.
\end{equation}
The following result is obvious from the assumptions and definitions.

\medskip

\begin{lem}
Let $A$ satisfy the decay condition in Assumption~\ref{assump}\textup{(\ref{assump2})}. Then we have
\begin{equation}
\We\in\cB(H^{1,s},H^{-1,s+\beta_2})\quad \text{for all $s\in\bR$}.
\end{equation}
\end{lem}

\begin{lem}\label{lemma23}
We have
\begin{equation}\label{vanish}
\ip{\psi_0}{(P_jA_j+A_jP_j)\psi_0}=0,\quad j=1,2,3,
\end{equation}
hence
\begin{equation}
\ip{\psi_0}{\We\psi_0}=\eps\ip{\psi_0}{\abs{A}^2\psi_0}.
\end{equation}
\end{lem}
\begin{proof}
The result \eqref{vanish} follows from Assumption~\ref{assump}(\ref{assump3}), since we take $\psi_0$ to be real-valued.
\end{proof}

We introduce the notation
\begin{equation}\label{YZdef}
\We=\eps^{1/2}Y+\eps Z,\quad Y=-A\cdot P-P\cdot A,\quad Z=\abs{A}^2.
\end{equation}
together with 
\begin{equation}\label{b-def}
b=\ip{\psi_0}{\abs{A}^2\psi_0}=\ip{\psi_0}{Z\psi_0}.
\end{equation}

In \cite{JKmag} a factorization of $\We$ is an essential tool. We recall its definition.
Take as the intermediate space $\cK=L^2(\bR^3;\bC^9)$.
First factor $A$ as follows:
\begin{equation}
A_j(s)=a_j(x)b_j(x),\quad a_j(x)=\abs{A_j(x)}^{1/2},
\quad b_j(x)=\sign(A_j(x))a_j(x),\quad j=1,2,3.
\end{equation}

We define
\begin{equation}
\we\colon H^{1,-\beta_2/2}\to \cK
\end{equation}
as a $9\times 1$ operator matrix. The entries are (omitting the column index)
\begin{alignat*}{2}
(\we)_j&=\eps^{1/2}A_j,& j&=1,2,3,\\
(\we)_j&=\eps^{1/4}a_{j-3},& j&=4,5,6,\\
(\we)_j&=\eps^{1/4}b_{j-6}P_{j-6},\quad& j&=7,8,9.
\end{alignat*}
Define the following self-adjoint and unitary operator on $\cK$:
\begin{equation*}
\cU=\begin{bmatrix}
\bone_3 & 0 & 0\\
0 & 0& -\bone_3\\
0 & -\bone_3  & 0
\end{bmatrix},
\end{equation*}
where the entries are $3\times3$ operator matrices, i.e.~the identity and the zero operator matrix.

Consider $\We$ as a bounded operator from $H^{1,-\beta_2/2}$ to
$H^{-1,\beta_2/2}$. Then we have the factorization
\begin{equation}\label{factor}
\We=\wes\cU\we.
\end{equation}

\begin{lem}\label{lemma24}
Let Assumption~\ref{assump}\textup{(\ref{assump2})} hold. Then for $0\leq s\leq\beta_2/2$ we have $\we\in\cB(H^{1,-s},\cK)$ and $\wes\in\cB(\cK,H^{-1,s})$, with the estimates
\begin{equation}
\norm{\we}_{\cB(H^{1,-s},\cK)}\leq C\eps^{1/4}
\quad\text{and}\quad
\norm{\wes}_{\cB(\cK,H^{-1,s})}\leq C\eps^{1/4}.
\end{equation}
\end{lem}

\section{Resolvent expansions}
We will need the asymptotic expansion of the resolvent $(H_0-z)^{-1}$ as $z\to0$ in the topology of $\cB(H^{-1,s},H^{1,-s'})$. We have the following result. As in \cite{JKmag} we change
to $\kappa$, where for $z\in\bC\setminus[0,\infty)$ we take $\kappa=-i\sqrt{z}$, with
$\im \sqrt{z}>0$, such that $z=-\kappa^2$.

\medskip

\begin{prop}\label{prop31}
Let $N\geq0$. Assume that $V$ satisfies Assumptions~\ref{assump}\textup{(\ref{assump1})} and~\ref{assump}\textup{(\ref{assump3})} for some $\beta_1>2N+9$. Then we have
\begin{equation}\label{eq31}
(H_0+\kappa^2)^{-1}=\frac{1}{\kappa^2}G_{-2}+\frac{1}{\kappa}G_{-1}
+\sum_{j=0}^N\kappa^jG_j+\kappa^{N+1}\widetilde{G}_N(\kappa)
\quad \text{as $\kappa\to0$}
\end{equation}
in the topology of $\cB(H^{-1,s},H^{1,-s'})$ for $s,s'>N+\frac92$.
Here $G_{-2}=P_0=\ket{\psi_0}\bra{\psi_0}$ is the eigenprojection onto eigenvalue $0$ and $G_{-1}=c_0P_0$, where
\begin{equation}\label{eq32}
c_0=\frac{1}{12\pi}(X_1^2+X_2^2+X_3^2), \quad
X_j=\int_{\bR^3}\psi_0(x)V(x)x_jdx,
\quad j=1,2,3.
\end{equation}
The error term satisfies $\norm{\widetilde{G}_N(\kappa)}_{\cB(H^{-1,s},H^{1,-s'})}\leq C$ for
$\abs{\kappa}\leq \delta$, for some $\delta>0$. The constant $C$ depends on $s$, $s'$, and $\delta$.
\end{prop}
The proof of this proposition is a variant of the proofs in~\cite{JK,J80,M,JN}. We omit the details.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{The SLFG formula}
\label{sec-SLFG}
We need the SLFG formula from~\cite{JN} for the Hamiltonians considered here.
As above $P_0$ is the eigenprojection for eigenvalue $0$ for $H_0$. Let $Q_0=I-P_0$. We write $\cH=P_0\cH\oplus Q_0\cH$. Let $R_{\eps}(z)=(H_{\eps}-zI)^{-1}$. We denote the resolvent of $Q_0H_{\eps}Q_0$ in the space $Q_0\cH$ by $R_{0,\eps}(z)$. i.e. $R_{0,\eps}(z)=(Q_0H_{\eps}Q_0-zQ_0)^{-1}$.

Then we have the matrix representation
\begin{equation*}
R_{\eps}(z)=\begin{bmatrix}
\Reff(z) & R_{12}(z)\\ R_{21}(z) & R_{22}(z)
\end{bmatrix},
\end{equation*}
where
\begin{align*}
\Reff(z)&=\bigl(P_0H_{\eps}P_0-P_0\We Q_0R_{0,\eps}(z)Q_0\We P_0-zP_0\bigr)^{-1},
\\
R_{12}(z)&=-\Reff(z)P_0\We Q_0 R_{0,\eps}(z),
\\
R_{21}(z)&=-R_{0,\eps}(z)Q_0\We P_0\Reff(z),
\\
R_{22}(z)&=R_{0,\eps}(z)+R_{0,\eps}(z)
Q_0\We P_0\Reff(z)P_0\We Q_0R_{0,\eps}(z).
\end{align*}
With a slight abuse of notation we write
\begin{equation*}
\Reff(z)=(\Heff(z)-zP_0)^{-1}\quad\text{on $P_0\cH$}
\end{equation*}
and
\begin{equation} \label{eq-eff}
P_0(\Heff(z)-z)P_0=F(z,\eps)P_0,
\end{equation}
where
\begin{align}
F(z,\eps)&=\ip{\psi_0}{\We\psi_0}-z
-\ip{\psi_0}{\We Q_0 R_{0,\eps}(z)Q_0\We\psi_0}
\notag\\
&=\eps\ip{\psi_0}{\abs{A}^2\psi_0}-z
-\ip{\psi_0}{\We Q_0 R_{0,\eps}(z)Q_0\We\psi_0}
\end{align}
by Lemma~\ref{lemma23}.

We note that on $Q_0\cH$
\begin{equation*}
Q_0(H_0-z)^{-1}Q_0=\bigl(Q_0H_0Q_0-zQ_0\bigr)^{-1}
\end{equation*}
and introduce the notation
\begin{equation}
G_{\eps}(z)=\we Q_0(H_0-z)^{-1}Q_0\wes.
\end{equation}
Using the factorization \eqref{factor} and the mapping properties of $\we$ and $\wes$ we can derive the analogue of \cite[Equation (2.14)]{JN}:
\begin{align}
F(z,\eps)&=\eps\ip{\psi_0}{\abs{A}^2\psi_0}-z
%\notag\\
%&\quad
-\ip{\psi_0}{\wes\cU\{G_{\eps}(z)-G_{\eps}(z)
[\cU+G_{\eps}(z)]^{-1}G_{\eps}(z)\}\cU\we\psi_0}.
\label{eq43}
\end{align}
We omit the details of this derivation.

%To use this result we need the following assumption.
%\begin{assumption}
%Assume that $b=\ip{\psi_0}{\abs{A}^2\psi_0}>0$.
%\end{assumption}

We would now like to use the resolvent expansion from Proposition~\ref{prop31} to get an expansion of $G_{\eps}(z)$. There is an important point to be made. We cannot just apply $Q_0$ on both sides of the expansion of $(H_0-z)^{-1}$. The reason is that the eigenfunction $\psi_0$ may not decay fast enough to ensure that
$Q_0H^{-1,s}\subset H^{-1,s}$ for the $s$ so large that the expansion holds.
Instead, we observe that on the space $\cH$ we have
\begin{equation*}
Q_0(H_0-z)^{-1}Q_0=(H_0-z)^{-1}+\frac{1}{z}P_0.
\end{equation*}
The right hand side has an asymptotic expansion in the topology $\cB(H^{-1,s},H^{1,-s})$, see Proposition~\ref{prop31}. We now switch to the $\kappa$ variable instead of the $z$ variable.

\medskip

\begin{prop}\label{prop42}
Let $N\geq0$.
Let Assumption~\ref{assump} hold with $\beta_j>2N+9$, $j=1,2$. Then the following asymptotic expansion holds in the topology of $\cB(\cK)$, as $\kappa\to0$.
\begin{equation}\label{Geps-expand}
G_{\eps}(\kappa)=\sum_{j=-1}^N\kappa^j\widetilde{G}_{\eps,j}+\kappa^{N+1}\widehat{E}_N(\kappa,\eps).
\end{equation}
We have for $j=-1,0,\ldots,N$
\begin{equation}\label{Geps-expand-c}
\widetilde{G}_{\eps,j}=\we G_j\wes,\quad \widehat{E}_N(\kappa,\eps)=\we\widetilde{G}_N(\kappa)\wes.
\end{equation}
with
\begin{equation}
\norm{\widetilde{G}_{\eps,j}}_{\cB(\cK)}\leq C_1\eps^{1/2},
\quad \norm{\widehat{E}_N(\kappa,\eps)}_{\cB(\cK)}\leq C_2\eps^{1/2},
\end{equation}
and $C_2$ independent of $\kappa$ for $\abs{\kappa}\leq1$.
\end{prop}
\begin{proof}
The results follow from Proposition~\ref{prop31} and Lemma~\ref{lemma24}. % We omit the details.
\end{proof}

\begin{assumption}\label{nu-assump}
Let $\nu\geq-1$ be an odd integer. Assume that $G_j=0$ for $j=-1,1,\ldots,\nu-2$.
Assume that $\ip{\psi_0}{YG_{\nu}Y\psi_0}\neq0$.
\end{assumption}

\begin{assumption}\label{pos-assump}
If $\nu=-1$, assume that
\begin{equation}\label{assump-b}
 b-2\ip{\psi_0}{YG_0Y\psi_0}>0.
\end{equation}
If $\nu\geq1$, assume that
\begin{equation}\label{assump-bb}
 b-\ip{\psi_0}{YG_0Y\psi_0}>0.
\end{equation}
\end{assumption}

Define
\begin{equation}\label{tb-def}
\tb=b-\ip{\psi_0}{YG_0Y\psi_0},
\end{equation}
with  $b$ given by \eqref{b-def}, and note that $\tb>0$.

We define
\begin{equation}\label{r-def}
r(\nu)=\begin{cases}
2, & \nu=-1,\\
4, & \nu=1,\\
\tfrac12(\nu+8), & \nu\geq3.
\end{cases}
\end{equation}
 Define
\begin{equation}\label{Deps}
\Deps=\{z=x+i\eta\mid
\tfrac12\tb\eps<x<\tfrac32\tb\eps,\;0<\abs{\eta}<(\tb\eps)^{r(\nu)}\}.
\end{equation}
We freely mix the notation $z$ and $\kappa$ for points in $\Deps$. Note that if $z\in \Deps$ then we have a lower bound $\abs{\kappa}=\abs{z}^{1/2}>C\eps^{1/2}$
where $C$ depends only on the parameter entering into the definition of $\Deps$. The power $r(\nu)$ is chosen to ensure that for $z\in\Deps$ the imaginary part $\im z$ can be absorbed in the error term in the expansion of $F(z,\eps)$, cf. Lemmas~\ref{lemma44}--\ref{lemma46}.

We write in statements below `for sufficiently small $\eps$' instead of the more precise formulation `there exists $\eps_0$ such that the statement holds for all $\eps$, $0<\eps<\eps_0$.'

We now state the essential decomposition and expansion results. Let
\begin{equation}\label{eq427}
\beta_{\nu}=\ip{\psi_0}{YG_{\nu}Y\psi_0}, \qquad \nu =1,3,5...
\end{equation}

\medskip

\begin{lem}[Case $\nu=-1$]
\label{lemma44}
Let Assumption~\ref{nu-assump} hold for $\nu=-1$.
For $\eps$ sufficiently small and $z\in \Deps$ we have
\begin{equation}\label{F-nu=-1}
F(z,\eps)=H(z,\eps)+r(z,\eps)
\end{equation}
with
\begin{equation}
\sup_{\substack{0<\eps<\eps_0\\ z\in\Deps}}\eps^{-2}\abs{r(z,\eps)}<\infty.
\end{equation}
and
\begin{equation}\label{eq412}
H(z,\eps)=\eps\tb-z-\eps^{3/2}\alpha_{-1}-
\kappa^{-1}\eps^2\beta_{-1},
\end{equation}
Here $\tb$ is defined in \eqref{tb-def} and $\alpha_{-1}, \beta_{-1}$ are real numbers given by
\begin{equation}\label{beta-1-def}
\alpha_{-1}=\ip{\psi_0}{(YG_0Z+ZG_0Y-YG_0YG_0Y)\psi_0}
\end{equation}
and
\begin{equation}\label{g-1-def}
\beta_{-1}=bc_0\bigl(b-2\ip{\psi_0}{YG_0Y\psi_0}\bigr).
\end{equation}
By Assumption \ref{pos-assump} we have $\beta_{-1}>0$.
\end{lem}

\medskip

\begin{lem}[Case $\nu=1$]\label{lemma45}
For $\eps$ sufficiently small and $z\in \Deps$ we have
\begin{equation}\label{F-nu=1}
F(z,\eps)=H(z,\eps)+r(z,\eps)
\end{equation}
with
\begin{equation}
\sup_{\substack{0<\eps<\eps_0\\ z\in\Deps}}\eps^{-4}\, \abs{r(z,\eps)}<\infty.
\end{equation}
We have
\begin{equation}\label{eq417}
H(z,\eps)=\eps(\tb-\kappa^2\alpha_2-\kappa^4\alpha_4)-z-\eps\kappa (\beta_1+\kappa^2\alpha_3+\kappa^4\alpha_5)
+\eps^{3/2}f(z,\eps)+\eps^{3/2}\kappa g(z,\eps).
\end{equation}
Here $\tb$ are $\beta_1$ are given by \eqref{tb-def} and \eqref{eq427}, and
\begin{align}
%\beta_1 & =\ip{\psi_0}{YG_1Y\psi_0}  \label{beta-1}, \\[5pt]
\alpha_j & =\ip{\psi_0}{YG_jY\psi_0},\quad j=2,3,4,5,
\end{align}
and
$f(z,\eps)$ and $g(z,\eps)$ are polynomials in $z$ with real coefficients. We have the following expressions:
\begin{align}
f(z,\eps)&=c_{\frac32,0}+\kappa^2 c_{\frac32,2}+\kappa^4 c_{\frac32,4}
+\eps^{1/2}\bigl(c_{2,0} + \kappa^2 c_{2,2}\bigr)\notag\\
&\quad
+\eps\bigl(c_{\frac52,0}+\kappa^2 c_{\frac52,2}\bigr)
+\eps^{3/2}C_{3,0}+\eps^2 c_{\frac72,0}
\end{align}
and
\begin{equation}
g(z,\eps)=c_{\frac32,1}+\kappa^2c_{\frac32,3}
+\eps^{1/2}\bigl(c_{2,1}+\kappa^2c_{2,3}\bigr)
+\eps c_{\frac52,1}+\eps^{3/2}c_{3,1}.
\end{equation}
\end{lem}

\medskip

\begin{lem}[Case $\nu\geq3$]
\label{lemma46}
Let Assumption~\ref{nu-assump} hold for some $\nu\geq3$. Then we can write
\begin{equation}
F(z,\eps)=H(z,\eps)+r(z,\eps),
\end{equation}
where
\begin{equation}
\sup_{\substack{0<\eps<\eps_0\\ z\in\Deps}}\eps^{-(\nu+8)/2}\abs{r(z,\eps)}<\infty.
\end{equation}
We have
\begin{align}
H(z,\eps)&=\eps\tb-z-\eps\kappa^{\nu}\bigl[\beta_{\nu}+\kappa^2a_2(\eps)
+\kappa^4a_4(\eps)\bigr] +\eps^{3/2}f(z,\eps) + \eps\kappa^2g(z,\eps),
\label{eq425}
\end{align}
with $\beta_\nu$ given by \eqref{eq427}.
The terms $a_2(\eps)$ and $a_4(\eps)$ are polynomials in $\eps^{1/2}$ with real coefficients. Furthermore, $f(z,\eps)$ and $g(z,\eps)$ are polynomials in $\kappa^2$ and $\eps^{1/2}$ with real coefficients. We have
\begin{equation*}
f(z,\eps)=f_0+\cO(\eps^{1/2})\quad\text{and}\quad
g(z,\eps)=g_0+\cO(\eps^{1/2}),
\end{equation*}
for $z\in\Deps$ and $\eps\to0$.
\end{lem}

To prove these lemmas we need the following result.

\medskip

\begin{lem}\label{lemma47}
For sufficiently small $\eps$ and $z\in \Deps$ the operator
$\cU+G_{\eps}(\kappa)$ is invertible in $\cB(\cK)$ with
$\norm{[\cU+G_{\eps}(\kappa)]^{-1}}$
uniformly bounded in $z$ and $\eps$. An asymptotic expansion of $[\cU+G_{\eps}(\kappa)]^{-1}$ can be obtained using the finite Neumann expansion.
\end{lem}
\begin{proof}
We give the proof in the case $\nu=-1$. The arguments in the other cases
are straightforward.
Write $\cU+G_{\eps}(\kappa)=\cU(I+\cU G_{\eps}(\kappa))$ and define
$X=\kappa^{-1}\cU\we G_{-1}\wes$. Using Proposition~\ref{prop42} we have
\begin{equation*}
\cU G_{\eps}(\kappa)=X+\cO(\eps^{1/2}).
\end{equation*}
For $z\in \Deps$ we only get a uniform bound on $X$, but we need decay in $\eps$. Using  Lemma~\ref{lemma23} we obtain
\begin{equation*}
X^2=\kappa^{-2}\cU\we G_{-1}\We G_{-1}\wes=
\kappa^{-2}c_0^2\cU\we P_0(\eps^{1/2}Y+\eps Z) P_0\wes
=\eps\kappa^{-2}c_0^2\cU\we P_0ZP_0\wes,
\end{equation*}
where $c_0$ is given in \eqref{eq32}. It follows that $\norm{X^2}_{\cB(\cK)}\leq C\eps^{1/2}$ for $z\in \Deps$.

Thus for $\eps$ sufficiently small $(I-X^2)^{-1}$ exists and can be computed using the Neumann series. Then we have
\begin{equation*}
(I+X)^{-1}=(I-X)(I-X^2)^{-1}.
\end{equation*}
This implies $(I+X)^{-1}=I-X+\cO(\eps^{1/2})$. Thus we have
\begin{equation*}
(I+\cU G_{\eps}(\kappa))^{-1}=(I+X)^{-1}\bigl(I+\cO(\eps^{1/2})(I+X)^{-1}\bigr)^{-1}.
\end{equation*}
The results follow from this expression and the Neumann series.
\end{proof}

\begin{proof}[\bf Proof of Lemma~\ref{lemma44}]
The starting point is \eqref{eq43}. The last term in \eqref{eq43} is split as follows
\begin{align}
&\ip{\psi_0}{\wes\cU\{G_{\eps}(z)-G_{\eps}(z)
[\cU+G_{\eps}(z)]^{-1}G_{\eps}(z)\}\cU\we\psi_0}
=\ip{\psi_0}{\wes\cU G_{\eps}(z)\cU \we\psi_0}
\label{term1}
\\[6pt]
&\qquad\qquad\qquad\qquad\qquad-\ip{\psi_0}{\wes\cU G_{\eps}(z)
[\cU+G_{\eps}(z)]^{-1}G_{\eps}(z)\cU\we\psi_0}
\label{term2}
\end{align}

We start by analyzing the term on the right hand side of \eqref{term1}. Rewrite, using Proposition~\ref{prop42}, as follows:
\begin{align}\label{expand_first}
\ip{\psi_0}{\wes\cU G_{\eps}(z)\cU \we\psi_0}
&=\sum_{j=-1}^N\kappa^j\ip{\psi_0}{\We G_j \We\psi_0}
\kappa^{N+1}\ip{\psi_0}{\We\widetilde{G}_N(\kappa,\eps)\We\psi_0}.
\end{align}

We can rewrite the $j=-1$ term using \eqref{YZdef}:
\begin{align}
\ip{\psi_0}{\We G_{-1}\We\psi_0}
&=\ip{\psi_0}{(\eps^{1/2}Y+\eps Z) G_{-1}(\eps^{1/2}Y+\eps Z)\psi_0}
\notag\\[5pt]
&=\eps\ip{\psi_0}{YG_{-1}Y\psi_0}+
\eps^{3/2}\ip{\psi_0}{(ZG_{-1}Y+YG_{-1}Z)\psi_0}
\notag\\[5pt]
&\quad+\eps^2\ip{\psi_0}{ZG_{-1}Z\psi_0}.
\label{last}
\end{align}

Combining $G_{-1}=c_0P_0$ with Lemma~\ref{lemma23} we find that the only nonzero term on the right hand side is the last one. Thus we have shown that
\begin{equation}
\ip{\psi_0}{\We G_{-1}\We\psi_0}=c_0b^2\eps^2,
\end{equation}
using $b=\ip{\psi_0}{\abs{A}^2\psi_0}=\ip{\psi_0}{Z\psi_0}$ and $c_0$, defined in \eqref{eq32}.
Taking into account the $\kappa^{-1}$ factor and $z\in \Deps$ it follows that the $j=-1$ contribution to the sum in~\eqref{expand_first} is of order $\eps^{3/2}$.

Next we analyze the $j=0$ term in the expansion~\eqref{expand_first}. Proceeding as above we get
\begin{align}
\ip{\psi_0}{\We G_{0}\We\psi_0}&=\eps\ip{\psi_0}{YG_0Y\psi_0}
+\eps^{3/2}\bigl(
\ip{\psi_0}{YG_0Z\psi_0} + \ip{\psi_0}{ZG_0Y\psi_0}\bigr)
\notag
\\
&\quad+\eps^2\ip{\psi_0}{ZG_0Z\psi_0}.
\end{align}
Thus we have a term of order $\eps$ which must be taken together with the first term in \eqref{eq43}.

The terms in \eqref{expand_first} with $1\leq j\leq N$ are of order $\eps^{(2+j)/2}$. The error term is of order $\eps^{(N+3)/2}$.

Next we use Lemma~\ref{lemma47} to deal with the term \eqref{term2}.
We rewrite it as follows:
\begin{align*}
\ip{\psi_0}{\wes\cU G_{\eps}(z)
[\cU+G_{\eps}(z)]^{-1}G_{\eps}(z)\cU\we\psi_0}&=
\ip{\psi_0}{\wes\cU G_{\eps}(z)\cU G_{\eps}(z)\cU\we\psi_0}
\\
&\quad-\ip{\psi_0}{\wes\cU G_{\eps}(z)\cU G_{\eps}(z)\cU G_{\eps}(z)\cU\we\psi_0}
\\
&\quad+\mathcal{O}(\eps^2)
\end{align*}
Consider the first term on the right hand side and apply the expansion in Proposition~\ref{prop42}.
\begin{align*}
\ip{\psi_0}{\wes\cU G_{\eps}(z)\cU G_{\eps}(z)\cU\we\psi_0}
&=
\kappa^{-2}\ip{\psi_0}{\We G_{-1}\We G_{-1}\We\psi_0}
\\
&\ +\kappa^{-1}\bigl(\ip{\psi_0}{\We G_{-1}\We G_{0}\We\psi_0}
%\\
%&\qquad
+\ip{\psi_0}{\We G_{0}\We G_{-1}\We\psi_0}\bigr)
\\
&\ +\kappa^0A_0(\eps)  +\cO(\kappa\eps^2),
\end{align*}
where
\begin{equation*}
A_0(\eps)= \ip{\psi_0}{\We G_{0}\We G_{0}\We\psi_0}+\ip{\psi_0}{\We G_{-1}\We G_1\We\psi_0}+\ip{\psi_0}{\We G_{1}\We G_{-1}\We\psi_0}\, .
\end{equation*}
The $\kappa^{-2}$ term is simplified using \eqref{YZdef} and Lemma~\ref{lemma23}.
The result is (recall $b=\ip{\psi_0}{Z\psi_0}$)
\begin{equation*}
\kappa^{-2}\ip{\psi_0}{\We G_{-1}\We G_{-1}\We\psi_0}=
\kappa^{-2}c_0^2b^3\eps^3.
\end{equation*}
In a similar manner the $\kappa^{-1}$ terms are rewritten as
\begin{align*}
\kappa^{-1}c_0b\bigl[2\eps^2\ip{\psi_0}{YG_0Y\psi_0}
+\eps^{5/2}\bigl(\ip{\psi_0}{YG_0Z\psi_0}&+\ip{\psi_0}{ZG_0Y\psi_0}\bigr)
+2\eps^3\ip{\psi_0}{ZG_0Z\psi_0}\bigr].
\end{align*}
Finally we also rewrite the $\kappa^{0}$ term as follows;
\begin{align*}
A_0(\eps) &=
\eps^{3/2} \ip{\psi_0}{YG_0YG_0Y\psi_0}
+ \eps^2\bigl(\ip{\psi_0}{ZG_0YG_0Y\psi_0}
+\ip{\psi_0}{YG_0ZG_0Y\psi_0}
\\
&\ +\ip{\psi_0}{YG_0YG_0Z\psi_0}\bigr) +\eps^{5/2}\bigl(
\ip{\psi_0}{ZG_0ZG_0Y\psi_0}
+\ip{\psi_0}{YG_0ZG_0Z\psi_0}
\\
&\ +\ip{\psi_0}{ZG_0YG_0Z\psi_0}\bigr)+\eps^3\ip{\psi_0}{ZG_0ZG_0Z\psi_0}
\\
&\ +2\eps c_0 b\bigl[
\eps\ip{\psi_0}{YG_1Y\psi_0} + \eps^{3/2}(\ip{\psi_0}{YG_1Z\psi_0}
+\ip{\psi_0}{ZG_1Y\psi_0}
\bigr)
+\eps^2\ip{\psi_0}{ZG_1Z\psi_0}
\bigr].
\end{align*}
The error term $\cO(\kappa\eps^2)$ follows by analyzing the structure of subsequent terms in the expansion.

The results in the lemma are a reorganized version of the above fairly detailed computations.
\end{proof}

\begin{proof}[\bf Proof of Lemma~\ref{lemma45}]
The proof is very similar to the proof of Lemma~\ref{lemma44}.
We omit the details.
\end{proof}
\begin{proof}[\bf Proof of Lemma~\ref{lemma46}]
We assume that $\nu\geq3$ is an odd integer such that Assumption~\ref{nu-assump} holds. Let us analyze how to get the right decomposition. The goal is to get an error term of order $\eps^{(\nu+8)/2}$.

We write
\begin{equation}
F(z,\eps)=\eps b-z+\Esf_1+\Esf_2,
\end{equation}
where
\begin{equation}
\Esf_1=-\ip{\psi_0}{\wes\cU G_{\eps}(z)\cU\we\psi_0}
\end{equation}
and
\begin{equation}
\Esf_2=\ip{\psi_0}{\wes\cU G_{\eps}(z)
[\cU+G_{\eps}(z)]^{-1}G_{\eps}(z)\cU\we\psi_0}.
\end{equation}
We have
\begin{equation} \label{E1-exp}
\Esf_1=-\ip{\psi_0}{\We\Bigl(
\sum_{j=0}^N\kappa^jG_j\Bigr)\We\psi_0}
+\kappa^{N+1}\ip{\psi_0}{\We \widetilde{G}_N(\kappa)\We\psi_0}
\end{equation}
The second term should go into the error term $r(z,\eps)$. Each $\We$  contributes $\eps^{1/2}$ and $\kappa\sim\eps^{1/2}$. Thus
\begin{equation*}
\kappa^{N+1}\eps\sim \eps^{1+(N+1)/2}.
\end{equation*}
Then $1+(N+1)/2=(\nu+8)/2$ shows that we must take $N=\nu+5$ in \eqref{E1-exp}.

Due to the assumption on the $G_j$ we split the sum above as
\begin{equation}\label{even-odd1}
\sum_{j=0}^N\kappa^jG_j=\sum_{k=0}^{(\nu+5)/2}\kappa^{2k}G_{2k}
+\kappa\sum_{l=(\nu-1)/2}^{(\nu+3)/2}\kappa^{2l}G_{2l+1}.
\end{equation}
Due to the two $\We$ factors we get the powers $\eps$, $\eps^{3/2}$ and
$\eps^2$. Multiplying out terms and reordering them we find
\begin{align*}
\kappa\!\!\!\sum_{l=(\nu-1)/2}^{(\nu+3)/2}\kappa^{2l} \ip{\psi_0}{\We\, G_{2l+1}\,  \We\psi_0}
&= \kappa^{\nu}\eps\bigl[
(c_{1,\nu}+\kappa^2c_{1,\nu+2}+\kappa^4c_{1,\nu+4})
\\
&
\quad +\eps^{\frac12}(c_{\frac32,\nu}+\kappa^2c_{\frac32,\nu+2}
+\kappa^4c_{\frac32,\nu+4})+\eps(c_{2,\nu}+\kappa^2c_{2,\nu+2})
\bigr]
\end{align*}
where one term was absorbed in the error term. We use the notation $c_{\alpha,\beta}$ for the coefficient to $\eps^{\alpha}\kappa^{\beta}$ in the expansion.

A similar computation for the even terms in \eqref{even-odd1} yields the following contribution:
\begin{equation*}
\sum_{k=0}^{(\nu+5)/2}\kappa^{2k} \ip{\psi_0}{\We\, G_{2k}\,  \We\psi_0} =
\eps\bigl[
\sum_{k=0}^{(\nu+3)/2}\kappa^{2k}\bigl(c_{1,2k}+
\eps^{1/2}c_{\frac32,2k}+\eps c_{2,2k}\bigr)
+\kappa^{\nu+5}c_{1,\nu+5}
\bigr].
\end{equation*}	
This finishes our analysis of $\Esf_1$.

Next we need to analyze the contributions from $\Esf_2$. For the term
$[\cU+G_{\eps}(z)]^{-1}$ we use Lemma~\ref{lemma47}. We omit the details of the computation.  We get a contribution with the following structure:

\begin{equation*}
\Esf_2 = \sum_{j=2}^{\nu+6}\, \sum_{k=j+1}^{\min\{2(j+1),\nu+7\}}
 \Biggl(
\sum_{\substack{l=0\\ \text{$l$ even}}}^{\nu+7-k}\eps^{k/2}\kappa^l d_{j,k,l}
+\sum_{\substack{l=\nu\\ \text{$l$ odd}}}^{\nu+7-k}
\eps^{k/2}\kappa^l d_{j,k,l}
\Biggr) + \mathcal{O}(\eps^{(\nu+8)/2})   \, .
\end{equation*}
For the sum over $l$ odd we use the standard convention
that $\sum_{l=m}^n\cdots=0$ for $n<m$.

Collecting and rearranging the terms we get the results stated in the lemma.
\end{proof}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Main results}
\label{sec-main}
In this section we use the results from Lemmas~\ref{lemma44}--\ref{lemma46} to prove the main results. We start with the analysis of the term $H(z,\eps)$.
%We assume $\nu\geq3$.
 It is clear that the limits
\begin{equation*}
H_{\pm}(x,\eps)=\lim_{\eta\downarrow0}H(x\pm i\eta,\eps)
\end{equation*}
exist. We have a decomposition in real and imaginary parts that can be written as
\begin{equation*}
H_{\pm}(x,\eps)=R(x,\eps)\pm i I(x,\eps).
\end{equation*}
Recall that $H(z,\eps)$ is given, for $\nu=-1, \nu=1$ and $\nu\geq 3$,  by \eqref{eq412}, \eqref{eq417} and \eqref{eq425} respectively.
In either case we obtain
\begin{equation}
R(x,\eps)=\tb\eps -x +  \varphi_\nu(x,\eps),
% \eps^{3/2}f(x,\eps)+\eps x g(x,\eps).
\end{equation}
where $\varphi_\nu$ is a real-valued function of $x$ which satisfies
\begin{equation} \label{fi-nu}
\sup_{x\in (\frac12\tb\eps,\frac32\tb\eps)} |\varphi_\nu| = \mathcal{O}(\eps^{3/2}) , \qquad \sup_{x\in (\frac12\tb\eps,\frac32\tb\eps)} |\partial_x \varphi_\nu| = \mathcal{O}(\eps), \qquad \nu=-1,1,3\dots
\end{equation}

It follows that for $\eps$ sufficiently small we have
\begin{equation*}
R(\tfrac12\tb\eps)<0\quad\text{and}\quad R(\tfrac32\tb\eps)>0.
\end{equation*}
We also have
\begin{equation*}
\frac{dR}{dx}(x,\eps)=-1+\cO(\eps),\quad\eps\to0.
\end{equation*}
Thus there exists a unique $x_0(\eps)\in(\frac12\tb\eps,\frac32\tb\eps)$ such that
\begin{equation} \label{R-x0}
R(x_0(\eps),\eps)=0\quad\text{and}\quad x_0(\eps)=\tb\eps+\cO(\eps^{3/2}).
\end{equation}

Define
\begin{equation}
\Gamma(\eps)=-I(\xe,\eps).
\end{equation}
Due to Assumption~\ref{nu-assump} and the fact that
\begin{equation*}
I(x,\eps)=i^{\nu-1}x^{\nu/2}\eps\,  \beta_\nu +\cO(\eps^{2+\frac\nu2}) ,
\end{equation*}
see \eqref{eq427},
it follows from \cite[(3.15)]{JN} and Assumption~\ref{nu-assump} that $\Gamma(\eps)>0$. More precisely, the coefficients $G_j$ in the expansion \eqref{eq31} are selfadjoint, and $\im R(z)\geq0$ for $\im z\neq0$. Note also that due to the assumption
$\ip{\psi_0}{YG_{\nu}Y\psi_0}\neq0$ there exist $0<c_1<c_2$ such that
\begin{equation}
\begin{aligned}\label{Ibound}
c_1\, \eps^{3/2} & \leq \abs{I(x,\eps)}\leq c_2\, \eps^{3/2} \qquad\quad\ \nu =-1, 1\\[5pt]
c_1\, \eps^{(\nu+2)/2} & \leq \abs{I(x,\eps)}\leq c_2\,  \eps^{(\nu+2)/2} \qquad \nu\geq 3
\end{aligned}
\end{equation}
for all $x\in(\frac12\tb\eps,\frac32\tb\eps)$ and $\eps$ sufficiently small.

We need to fix a small interval $\Ie\subset(\frac12\tb\eps,\frac32\tb\eps)$.
Define
\begin{equation}\label{I-eps}
\Ie=\begin{cases}
[\xe-\frac14\tb\eps,\xe+\frac14\tb\eps], & \nu=-1,1,\\[8pt]
[\xe-\frac{\Gamma(\eps)}{\eps^2},\xe+\frac{\Gamma(\eps)}{\eps^2}], & \nu\geq3.
\end{cases}
\end{equation}
Note that for $\eps$ sufficiently small
$\Ie\subset(\frac12\tb\eps,\frac32\tb\eps)$.

Our goal is to obtain information on
\begin{equation}\label{Ageps}
A_{g_{\eps}}(t)=\ip{\psi_0}{e^{-itH_{\eps}}g_{\eps}(H_{\eps})\psi_0},
\end{equation}
where $g_{\eps}$ is the characteristic function of the interval $I_{\eps}$. Using Stone's formula and the SLFG formula we have
\begin{equation*}
A_{g_{\eps}}(t)=\lim_{\eta\downarrow0}\frac{1}{2\pi i}
\int_{I_{\eps}}e^{-itx}\Bigl(\frac{1}{F(x+i\eta,\eps)}-
\frac{1}{F(x-i\eta,\eps)}\Bigr)dx.
\end{equation*}

\begin{lem}\label{lemma51}
For sufficiently small $\eps$ we have
\begin{equation}
\biggl\lvert
A_{g_{\eps}}(t)-\frac{1}{2\pi i}\int_{I_{\eps}}e^{-itx}
\Bigl(\frac{1}{H_+(x,\eps)}-
\frac{1}{H_-(x,\eps)}\Bigr)dx
\biggr\rvert
\leq C\eps^{q(\nu)}
\end{equation}
with
\begin{equation}
q(\nu)=\begin{cases}
\frac12, & \nu=-1,\\[5pt]
\frac52, & \nu=1,\\[5pt]
3, & \nu\geq3.
\end{cases}
\end{equation}
\end{lem}

\begin{proof}
Assume first that  $\nu\geq3$. Lemma~\ref{lemma46} implies that
\begin{equation*}
\abs{F(z,\eps)}\geq\abs{H(z,\eps)}-C\eps^{(\nu+8)/2}
\end{equation*}
for $\eps$ sufficiently small and $z\in\Deps$. Then the estimate \eqref{Ibound} implies that we get
\begin{equation*}
\abs{F(z,\eps)}\geq\frac12\abs{H(z,\eps)}
\end{equation*}
for $\eps$ sufficiently small.
Then
\begin{equation}
\biggl\lvert
\frac{1}{2\pi i}
\int_{I_{\eps}}e^{-itx}\Bigl(\frac{1}{F(x\pm i\eta,\eps)}-
\frac{1}{H(x\pm i\eta,\eps)}\Bigr)dx
\biggr\rvert
\leq C\eps^{(\nu+8)/2}\int_{I_{\eps}}\frac{1}{\abs{H(x\pm i\eta,\eps)}^2}\ dx.
\end{equation}
We can take the limit $\eta\downarrow0$ to get
\begin{equation} \label{H-L-1}
\biggl\lvert\lim_{\eta\downarrow0}
\frac{1}{2\pi i}
\int_{I_{\eps}}e^{-itx}\Bigl(\frac{1}{F(x\pm i\eta,\eps)}-
\frac{1}{H(x\pm i\eta,\eps)}\Bigr)dx
\biggr\rvert
\leq C\eps^{(\nu+8)/2}\int_{I_{\eps}}\frac{dx}{\abs{H_{\pm}(x,\eps)}^2}
\end{equation}
We have
\begin{equation*}
\abs{H_{\pm}(x,\eps)}^2=\abs{R(x,\eps)}^2+\abs{I(x,\eps)}^2
=\abs{R(x,\eps)-R(x_0(\eps),\eps)}^2+\abs{I(x,\eps)}^2.
\end{equation*}
Using the expressions for $R(x,\eps)$ and $I(x,\eps)$, and the estimate \eqref{Ibound}, we get
\begin{equation}\label{Hpm-lower}
\abs{H_{\pm}(x,\eps)}^2\geq C\bigl(\abs{x-x_0(\eps)}^2+\eps^{\nu+2}\bigr)
\end{equation}
for some $C<1$, see Appendix \ref{sec-app} for details.
Then we get
\begin{align*}
\eps^{(\nu+8)/2}\int_{I_{\eps}}\frac{1}{\abs{H_{\pm}(x,\eps)}^2}\ dx&\leq
C\eps^3\int_{-\infty}^{\infty}
\frac{\eps^{(\nu+2)/2}}{(x-x_0(\eps))^2+\eps^{\nu+2}}\ dx
\\[5pt]
&
=C\eps^3\int_{-\infty}^{\infty}\frac{1}{1+x^2}\ dx=C\pi \eps^3.
\end{align*}
Now suppose that  $\nu\in\{-1,1\}$. We proceed as in the case $\nu\geq 3$ and using Lemmas \ref{lemma44} and \ref{lemma45}
we arrive at
\begin{equation} \label{H-L-2}
\biggl\lvert\lim_{\eta\downarrow0}
\frac{1}{2\pi i}
\int_{I_{\eps}}e^{-itx}\Bigl(\frac{1}{F(x\pm i\eta,\eps)}-
\frac{1}{H(x\pm i\eta,\eps)}\Bigr)dx
\biggr\rvert
\leq C\, \eps^{3+\nu}\int_{I_{\eps}}\frac{dx}{\abs{H_{\pm}(x,\eps)}^2}
\end{equation}
In view of \eqref{Ibound} estimate \eqref{Hpm-lower} is now replaced by
\begin{equation}\label{Hpm-lower-2}
\abs{H_{\pm}(x,\eps)}^2\geq C\bigl(\abs{x-x_0(\eps)}^2+\eps^3\bigr)
\end{equation}
From this point we follow the analysis of the case $\nu\geq 3$ line by line and get
$$
 \eps^{3+\nu}\int_{I_{\eps}}\frac{dx}{\abs{H_{\pm}(x,\eps)}^2} \, \leq\, C\,  \eps^{\frac 32+\nu}
$$
as required.
\end{proof}

Define
\begin{equation} \label{L-eq}
L_{\pm}(x,\eps)=-\bigl(x-x_0(\eps)\bigr)\pm iI(x_0(\eps),\eps).
\end{equation}

\begin{lem}\label{lemma52}
We have the following estimates:
\begin{equation}\label{eq511}
\Big\lvert
\frac{d}{dx}H(x,\eps)+1
\Big\rvert
\leq C\begin{cases}
\eps^{1/2}, & \nu=-1,1,\\
\eps, & \nu\geq3.
\end{cases}
\end{equation}
\begin{equation}\label{eq512}
\Big\lvert
\frac{d^2}{dx^2}H(x,\eps)
\Big\rvert
\leq C\begin{cases}
\eps^{-1/2}, & \nu=-1,1,\\
\eps^{1/2}, & \nu=3,\\
\eps, & \nu\geq5.
\end{cases}
\end{equation}
\end{lem}
\begin{proof}
Assume $\nu\geq3$. Then \eqref{eq425} in combination with \eqref{fi-nu} implies
\begin{align*}
\frac{d}{dx}H(x,\eps)&=-1+\frac{\eps(-i)^{\nu}}{2} \bigl[\beta_{\nu}\, \nu\, x^{\frac{\nu-2}{2}}
-a_2(\eps)(\nu+2)\, x^{\frac \nu2}+a_4(\eps)(\nu+4)x^{\frac{\nu+4}{2}}\bigr]
 +  \mathcal{O}(\eps) .
\end{align*}
This implies \eqref{eq511}.
The second derivative is estimated in the same manner.
The cases $\nu=-1$ and $\nu=1$ follow by using
\eqref{eq412} and \eqref{eq417}, respectively.
\end{proof}

\begin{lem}\label{lemma53}
For sufficiently small $\eps$ we have
\begin{equation}
\biggl\lvert
\int_{I_{\eps}}e^{-itx}\Bigl(\frac{1}{H_{\pm}(x,\eps)}-
\frac{1}{L_{\pm}(x,\eps)}\Bigr)dx
\biggr\rvert\leq C
\begin{cases}
\eps^{1/2}, & \nu=-1,1,\\[5pt]
\eps, & \nu\geq3.
\end{cases}
\end{equation}
\end{lem}
\begin{proof}
Let us start with the case $\nu=-1$. It follows from the definition of $L_{\pm}(x,\eps)$ that we have $H_{\pm}(\xe,\eps)-L_{\pm}(\xe,\eps)=0$.

Using Taylor's formula with expansion point $x_0(\eps)$ we get the two estimates
\begin{align}
\abs{H_{\pm}(x,\eps)-L_{\pm}(x,\eps)-
(1+\tfrac{d}{dx}H_{\pm}(x_0(\eps),\eps)(x-x_0(\eps))}&\leq C\eps^{-1/2}\abs{x-x_0(\eps)}^2
\label{est-1}\\
\abs{H{_\pm}(x,\eps)-L_{\pm}(x,\eps)}&\leq C\eps^{1/2}\abs{x-x_0(\eps)}
\label{est-2}
\end{align}

Rewrite as follows
\begin{equation}
\frac{1}{H_{\pm}(x,\eps)} - \frac{1}{L_{\pm}(x,\eps)}=
\frac{L_{\pm}(x,\eps)-H_{\pm}(x,\eps)}{L_{\pm}(x,\eps)^2}
+\frac{\bigl(L_{\pm}(x,\eps)-H_{\pm}(x,\eps)\bigr)^2}{L_{\pm}(x,\eps)^2H_{\pm}(x,\eps)}.
\label{est-3}
\end{equation}
Note that \eqref{Hpm-lower} holds in the $\nu=-1$ case with $\eps^{\nu+2}$ replaced by $\eps^3$. Then we have
\begin{equation*}
\abs{H_{\pm}(x,\eps)}\geq C\abs{x-x_0(\eps)}
\end{equation*}
Using this lower bound in combination with  \eqref{est-1}, \eqref{est-2}, and \eqref{est-3} it follows that
\begin{align}
\biggl\lvert
\int_{I_{\eps}}e^{-itx} \Bigl(\frac{1}{H_{\pm}(x,\eps)}-
\frac{1}{L_{\pm}(x,\eps)}\Bigr) dx
\biggr\rvert
&\leq
\abs{1+\tfrac{d}{dx}H_{\pm}(x_0(\eps))} \ \biggl\lvert
\int_{I_{\eps}}e^{-itx}\frac{x_0(\eps)-x}{L_{\pm}(x,\eps)^2}\ dx\biggr\rvert
\label{eq-a}\\[7pt]
&\quad
+C\eps^{-1/2}\int_{I_{\eps}}\frac{\abs{x-x_0(\eps)}^2}{\abs{L_{\pm}(x,\eps)}^2}\
dx
\label{eq-b}\\[7pt]
&\quad
+C\eps\int_{I_{\eps}}\frac{\abs{x-x_0(\eps)}}{\abs{L_{\pm}(x,\eps)}^2}\ dx.
\label{eq-c}
\end{align}

Recall the choice of $I_{\eps}$, see \eqref{I-eps}.   We have
\begin{equation*}
\frac{\abs{x-x_0(\eps)}^2}{\abs{L_{\pm}(x,\eps)}^2}\leq1 .
\end{equation*}
Since the length of $I_{\eps}$ equals $\frac12\tb\eps$, it follows that the term  \eqref{eq-b} is bounded by $C\eps^{1/2}$. The term \eqref{eq-c} is estimated using
\eqref{L-eq} as follows
\begin{equation*}
 \eps\int_{I_{\eps}}\frac{\abs{x-x_0(\eps)}}{\abs{L_{\pm}(x,\eps)}^2}\ dx \leq C
\eps\int_{I_{\eps}}\frac{1}{\sqrt{(x-\xe)^2+\eps^3}}\ dx
\leq C\eps\abs{\ln(\eps)}.
\end{equation*}
The integrand on the right hand side of \eqref{eq-a} is rewritten as
\begin{equation*}
\frac{x_0(\eps)-x}{L_{\pm}(x,\eps)^2}=\frac{1}{L_{\pm}(x,\eps)}
\pm\frac{i\Gamma(\eps)}{L_{\pm}(x,\eps)^2}.
\end{equation*}
Thus,
\begin{align*}
\biggl\lvert
\int_{I_{\eps}}e^{-itx}\frac{x_0(\eps)-x}{L_{\pm}(x,\eps)^2}\ dx\biggr\rvert
\leq \biggl\lvert
\int_{I_{\eps}}e^{-itx}\frac{1}{L_{\pm}(x,\eps)}dx
\biggr\rvert
+
\int_{I_{\eps}}\frac{\Gamma(\eps)}{\abs{L_{\pm}(x,\eps)}^2}\ dx
\end{align*}
Using the definition of $L_{\pm}(x,\eps)$ we find
\begin{equation*}
\int_{I(\eps)}\frac{\Gamma(\eps)}{\abs{L_{\pm}(x,\eps)}^2}\ dx
\leq \int_{\R}\frac{\Gamma(\eps)}{x^2+\Gamma(\eps)^2}\ dx=\pi.
\end{equation*}
Now define
\begin{equation} \label{length}
\ell(\eps)=\frac{\text{length}(I_{\eps})}{2}\, .
\end{equation}
Then
\begin{equation*}
\biggl\lvert
\int_{I_{\eps}}e^{-itx}\frac{1}{L_{\pm}(x,\eps)}dx
\biggr\rvert
=\biggl\lvert
\int_{-\ell(\eps)/\Gamma(\eps)}^{\ell(\eps)/\Gamma(\eps)}
e^{-it\Gamma(\eps)y}\frac{1}{y\mp i}dy
\biggr\rvert
\leq C.
\end{equation*}
Here, in the last step, we have used \cite[Eq.~(3.56)]{JN}, see also \cite[Lem.~B.1]{BK}.
Using the estimate \eqref{eq511}
the proof is finished in the case $\nu=-1$. The cases $\nu\geq1$ are treated in the same manner.
\end{proof}

Recall that  $\xe=\tb\eps+\cO(\eps^{3/2})$ and  $\Gamma(\eps)=-I(\xe,\eps)$.
We have the following asymptotics of $\Gamma(\eps)$.
\begin{equation}\label{gamma-asymp}
\Gamma(\eps)=
\begin{cases}
\tb^{-1/2}\beta_{-1}\, \eps^{3/2}+\cO(\eps^2), & \nu=-1,
\\[5pt]
\tb^{1/2}\beta_{1}\, \eps^{3/2}+\cO(\eps^{2}), & \nu=1,
\\[5pt]
-i^{\nu-1}\, \tb^{\nu/2}\, \eps^{(\nu+2)/2}\beta_{\nu}
+\cO(\eps^{(\nu+4)/2}), & \nu\geq3,
\end{cases}
\end{equation}
where the coefficients   $\beta_{-1}, \beta_1$, and $\beta_{\nu}$  are given by \eqref{g-1-def} and  \eqref{eq427}.

Let
\begin{equation}\label{p-eq}
p = \begin{cases}
\eps^{1/2}, & \nu=-1,1,\\[6pt]
\eps^2, & \nu\geq3.
\end{cases}
\end{equation}

\medskip

\begin{lem}\label{lemma54}
For $\eps$ sufficiently small we have
\begin{equation}\label{A-est}
\abs{A_{g_{\eps}}(t)-e^{-it(\xe-i\Gamma(\eps))}}
\leq C\, \eps^p\, .
\end{equation}
\end{lem}
\begin{proof}
From  Lemmas \ref{lemma51} and \ref{lemma53} it follows that
\begin{equation}
A_{g_{\eps}}(t) = \frac{1}{2\pi i}\int_{I_{\eps}}e^{-itx}\Bigl[
\frac{1}{L_+(x,\eps)}-\frac{1}{L_-(x,\eps)}
\Bigr]dx + \mathcal{O}(\eps^p),
\end{equation}
with an error term uniform in $t>0$.
We have
\begin{equation}\label{L-comp}
\frac{1}{2\pi i}\int_{I_{\eps}}e^{-itx}\Bigl[
\frac{1}{L_+(x,\eps)}-\frac{1}{L_-(x,\eps)}
\Bigr]dx
=\frac{1}{\pi}\int_{I_{\eps}}e^{-itx}
\frac{\Gamma(\eps)}{(x-x_0(\eps))^2+\Gamma(\eps)^2}dx.
\end{equation}
Since $\ell(\eps)/\Gamma(\eps)\to \infty$ as $\eps\to 0$, equations \eqref{length} and \eqref{gamma-asymp} imply
\begin{equation}\label{dif-comp}
\begin{aligned}
\Bigl\lvert\Bigl(\int_{\R}-\int_{I_{\eps}}\Bigr)
e^{-itx}
\frac{\Gamma(\eps)}{(x-x_0(\eps))^2+\Gamma(\eps)^2}dx\Bigr\rvert
&\leq
C \int_{\ell(\eps)}^{\infty}\frac{\Gamma(\eps)}{x^2+\Gamma(\eps)^2}dx \\[5pt]
&= C\int_{\ell(\eps)/\Gamma(\eps)}^\infty \frac{1}{1+t^2}\ dt
\leq  C\  \frac{\Gamma(\eps)}{\ell(\eps)} \\[5pt]
& \leq C  \eps^\nu  .
\end{aligned}
\end{equation}
By the residue theorem we have
\begin{equation*}
\frac{1}{\pi}\int_{\R}e^{-itx}
\frac{\Gamma(\eps)}{(x-x_0(\eps))^2+\Gamma(\eps)^2}dx
=e^{-it(x_0(\eps)-i\Gamma(\eps)}.
\end{equation*}
Estimate \eqref{A-est} thus follows from the definition of $I_{\eps}$, see \eqref{I-eps}, and \eqref{gamma-asymp}.
\end{proof}

We can now state our main result.

\begin{thm}\label{main}
Suppose that Assumptions~\ref{nu-assump} and~\ref{pos-assump} hold. Then for all $\lambda$ sufficiently small we have
\begin{equation}
\sup_{t>0} \abs{\ip{\psi_0}{\big[ e^{-itH_{\lambda}}-e^{-it(x_0(\lambda)-i\Gamma(\lambda))}\big] \psi_0} }
\leq C \lambda^{2p}\, ,
\end{equation}
where
\begin{equation}\label{gamma-lambda}
\Gamma(\lambda)=
\begin{cases}
\tb^{\, -\frac12}\beta_{-1}\, |\lambda|^3+\cO(\lambda^4), &\  \nu=-1,
\\[5pt]
\tb^{\, \frac 12}\beta_{1}\,  |\lambda|^3+\cO(\lambda^{4}), & \ \nu=1,
\\[5pt]
-i^{\nu-1}\, \tb^{\, \frac \nu2}\,  |\lambda|^{\nu+2}\, \beta_{\nu}
+\cO(|\lambda|^{\nu+4}), &\  \nu\geq3,
\end{cases}
\end{equation}
and
\begin{equation} \label{x0-asymp}
x_0(\lambda)=\tb \lambda^2+\cO(|\lambda|^{3}).
\end{equation}
Recall that $p$ is given by \eqref{p-eq}.
\end{thm}

\begin{proof}
Since the coefficients $b, \tb$ and $\beta_\nu, \nu=-1,2,3\dots,$ are even in $A$, the claim will follow
from the estimate
\begin{equation}\label{enough}
\sup_{t>0} \abs{\ip{\psi_0}{\big[ e^{-itH_{\eps}}-e^{-it(\xe-i\Gamma(\eps))}\big] \psi_0} }
\leq C\, \eps^p
\end{equation}
upon setting $\lambda=\sqrt{\eps}$.
With Lemma  \ref{lemma54} at hand, the proof of \eqref{enough} becomes a straightforward
variation on the proof of \cite[Theorem~3.7]{JN}.
Following the argument of Hunziker \cite{hun}, see also \cite{JN}, we first we apply equation
\eqref{A-est} with $t=0$ to get $|A_{g_{\eps}}(0)-1|   \lesssim  \eps^p$. Consequently, by \eqref{Ageps},
\begin{equation}  \label{eq-hunz-1}
\| \big(1- g_\eps (H_\eps )\big)^{\frac 12}\, \psi_0\|_2 \, \leq \, C  \eps^p.
\end{equation}
This implies
\begin{align} \label{eq-hunz-2}
\big | \LL \psi_0, \, e^{-itH_\eps}\, \psi_0\RR -A_{g_{\eps}}(t) \big | & = \big | \LL \big(1- g_\eps (H_\eps)\big)^{\frac 12}\,
\psi_0,\, e^{-it H_\eps}\big(1- g_\eps (H_\eps)  \big)^{\frac 12}\, \psi_0 \RR \big)  \nonumber \\
& \leq \| \big(1- g_\eps (H_\eps)  \big)^{\frac 12}\, \psi_0\|_2\, \leq\  C \eps^p.
\end{align}
Equation \eqref{enough} now follows from Lemma \ref{lemma54}. Since  $\Gamma(\eps)$ satisfies \eqref{gamma-asymp}, and
sine
$\xe=\tb\eps+\cO(\eps^{3/2})$, cf.~equation \eqref{R-x0}, this completes the proof.
\end{proof}

\begin{rem}
Notice that the leading term in the asymptotic expansion of $\Gamma$ is of the same order of magnitude for $\nu=-1$ and $\nu=1$. This is yet another 
important difference with respect to the resonances induced by an electric field, cf.~\cite[Thm.~3.7]{JN}.
\end{rem}

\begin{cor}\label{cor-unique}
Let the assumptions of Theorem \ref{main} be satisfied. Assume that there exist $\widehat x_0(\lambda)$ and $\widehat\Gamma(\lambda)$ such that
for all $\lambda$ small enough we have
$$
\sup_{t>0} \abs{\ip{\psi_0}{\big[ e^{-itH_{\lambda}}-e^{-it(\widehat x_0(\lambda)-i\widehat\Gamma(\lambda))}\big] \psi_0} }
\leq C\, \lambda^{2p}\, .
$$
Then
$$
\lim_{\lambda\to 0} \frac{\widehat\Gamma(\lambda)}{\Gamma(\lambda)} =\lim_{\lambda\to 0} \frac{\widehat x_0(\lambda)}{x_0(\lambda)} =1.
$$
\end{cor}
\begin{proof}
From \eqref{gamma-asymp} and from \cite[Prop.~1.1]{JN2} it follows that
$$
|\widehat\Gamma(\lambda)-\Gamma(\lambda)| + |\widehat x_0(\lambda)- x_0(\lambda)|  = \mathcal{O}(|\lambda|^{2p+2+|\nu|}) \qquad\lambda\to 0
$$
for all $\nu=-1,1,3\dots$. The claim thus follows from equations  \eqref{gamma-lambda} and \eqref{x0-asymp}.
\end{proof}

\appendix
\section{Comments on \eqref{Hpm-lower}}
\label{sec-app}
The bound
\begin{equation*}
\abs{H_{\pm}(x,\eps)}^2\geq C\bigl(\abs{x-x_0(\eps)}^2+\eps^{\nu+2}\bigr)
\end{equation*}
may require some explanation. We give the details here.
We have
\begin{equation*}
H_{\pm}(x,\eps)=R(x,\eps)-R(\xe,\eps)\pm i I(\xe,\eps).
\end{equation*}
Then
\begin{multline*}
R(x,\eps)-R(\xe,\eps)
\\
=\xe-x+\eps^{3/2}[f(x,\eps)-f(\xe,\eps)]
+\xe\eps g(\xe,\eps) - x\eps g(x,\eps).
\end{multline*}
$f(x,\eps)$ is a polynomial in $x$. Thus we have
\begin{equation*}
\eps^{3/2}[f(x,\eps)-f(\xe,\eps)]=\eps^{3/2}(x-\xe)[c+\cO(\eps)].
\end{equation*}
$g(x,\eps)$ is also a polynomial in $x$. Thus
\begin{align*}
x\eps g(\xe,\eps) - \xe\eps g(x,\eps)&=\eps(x-\xe)g(x,\eps)
+\eps\xe[g(x,\eps)-g(\xe,\eps)]
\\
&=\eps(x-\xe)[c+\cO(\eps)]
\end{align*}
Thus we have
\begin{equation*}
\abs{R(x,\eps)-R(\xe,\eps)}^2=\abs{x-\xe}^2(1+\cO(\eps^2))
\geq\tfrac12\abs{x-\xe}^2
\end{equation*}
for $\eps$ sufficiently small.
Since $\beta_{\nu}\neq0$, we have
\begin{equation*}
I(\xe,\eps)=i^{\nu-1}\eps(\xe)^{\nu/2}[\beta_{\nu}+\cO(\xe)],
\end{equation*}
and then
\begin{equation*}
\abs{I(\xe,\eps)}^2\geq C \eps^{\nu+2}.
\end{equation*}
The estimate \eqref{Hpm-lower} follows from these computations.

\end{document}